# A 26-Gram Butterfly-Inspired Robot Achieving Autonomous Tailless Flight

### v1

@@ Discussion
Calculated from available morphological data. ∗∗Estimated from geometry using published images. Reported as 3.81 in  (?), though calculated value gives $AR=3.29$ using the listed dimensions.
@@ Acknowledgments
We are grateful to Prof. G.C.H.E. de Croon, Prof. Maurizio Porfiri, and Prof. Lu Fang for their insightful comments and feedback on this manuscript. W.G. acknowledges support from the China Postdoctoral Science Foundation under Grant Number 2025M781650 and the Shuimu Scholar Program of Tsinghua University under Grant Number 2024SM224. This work was also supported by Xinchen Qihang Inc.
@@
Corresponding authors: Chao Gao, Guyue Zhou
Email: chao.gao@cantab.net, zhouguyue@air.tsinghua.edu.cn
@@
Wings were fabricated from Mylar film reinforced with carbon fiber spars (a 1.0 mm main spar, 0.8 mm leading edge and hindwing contour, and 0.5 mm longitudinal veins). The carbon reinforcement provided a high stiffness-to-weight ratio and enabled spatial stiffness gradients analogous to those found in butterfly wings. Alternative membranes were evaluated: TPU exhibited excessive compliance and poor adhesion, while silicone-coated polyamide (PA66) offered durability but imposed excessive mass. All tested membranes were flight-capable; however, PET was ultimately chosen for its balance of light weight, stiffness, tear resistance, adhesive compatibility, and optical transparency for motion capture. Wing fabrication involved laser-printing a venation template, overlaying the PET film, and bonding carbon rods using UV-curable adhesive and cyanoacrylate. Left-right wing pairs were mass-balanced within 0.05 g ($\leq 1\%$ of a single wing mass) to minimize asymmetries that could induce unbalanced aerodynamic forces or fuselage vibrations from uneven stroke timing or deformation  (?).
@@
where $\{\mathbf{P}_{i}\}_{i=0}^{n-1}$ are the spline control points determined by smoothing least-squares fitting, and $N_{i,3}(u)$ are the cubic B-spline basis functions defined by the Cox-deBoor recursion  (?) as
@@
Forewing images were collected primarily from published literature with documented clear venation patterns  (?), supplemented by photographs of museum-grade specimens available from open-access sources (Wikipedia). Specimens were selected based on strict visibility criteria requiring a fully unobstructed forewing (dorsal or ventral) view with distinct Dc and Cu1 veins, which define key geometric axes for comparative analysis. The dense and pigmented scales on butterfly wings often obscure vein boundaries, substantially constraining the number of usable samples. Consequently, the final dataset included 24 different specimens representing three families, namely Nymphalidae ($n=8$), Pieridae ($n=8$), and Papillionidae ($n=8$). For each specimen, the forewing image was first converted to grayscale and manually annotated with yellow solid lines indicating the Dc and Cu1 veins (fig. S2). The annotated regions were then segmented in HSV color space to extract the obtuse inter-vein Dc-Cu1 angle. Future morphological studies could expand statistical results by employing chemical bleaching or depigmentation techniques to enhance venation contrast and enable automated image segmentation.
@@
Wing performance was assessed using a custom flapping apparatus (Fig. 2D). Retroreflective stickers were selected for their low mass, which is the lightest available option despite still adding some weight. Increasing the number or size of markers could improve tracking accuracy but would also alter wing dynamics due to extra weight. Each wing variant underwent same flapping cycles while synchronized force and kinematic data were recorded at 222 Hz. The raw force signals were processed through a computational pipeline that first identified active flapping periods by detecting servo motor activation thresholds in the PWM command signals. Stroke reversal points were determined via second-order differentiation of positional data to precisely segment individual flapping cycles. Aerodynamic forces were isolated by subtracting inertial contributions using a phase-locked averaging approach. Specifically, a perforated counterpart was fabricated for each wing variant to approximate the original mass and inertial properties while substantially reducing aerodynamic loading. Forces measured under identical flapping conditions with these perforated wings were treated as inertial baselines and subtracted from the corresponding intact-wing measurements. For each wing configuration, 40 consecutive flapping cycles were selected from steady-state operation and processed with a fourth-order Butterworth low-pass filter (50 Hz cutoff) to remove high-frequency noise. The filtered force data were temporally normalized to the interval $[0,1]$ to enable cycle-wise comparison. Component-wise force signals ($F_{x}$, $F_{y}$, $F_{z}$) were then averaged and their variances computed to characterize aerodynamic force profiles across the normalized flapping period. Processed forces were further analyzed via polar coordinate transformation, where the angular coordinate represented normalized time ($\theta_{\mathrm{polar}}=2\pi t_{\mathrm{norm}}$) and the radial coordinate denoted force magnitude with an offset ensuring positivity (Fig. 2, F and G). Time-integrated forces were computed using trapezoidal numerical integration for both signed and absolute components, providing quantitative measures of net impulse and cumulative force production. This analysis enabled comprehensive comparison of force magnitude, temporal phasing, and generation patterns across different wing morphological variants.
@@
Wing deformation was quantified by projecting three-dimensional marker positions onto a plane orthogonal to the wing rotation axis, isolating bending motion from the overall flapping trajectory. Bending angles were defined as the angular deviation between proximal and distal segments, with maxima identified separately for downstroke and upstroke phases (Fig. 2, I and J). Hindwing kinematics were characterized by computing elevation angles relative to the wing hinge throughout each flapping cycle. Forewing kinematics were derived from markers placed along the anterior (leading edge) and posterior (trailing edge) margins of each wing section, enabling direct quantification of instantaneous wing orientation relative to the airflow. Discrete marker data were interpolated to generate continuous angle-time curves, enabling phase-resolved comparison of forewing and hindwing kinematics across wing variants.

### v2

@@ Results
Calculated from available morphological data. ∗∗Estimated from geometry using published images. Reported as 3.81 in [54], though calculated value gives $AR=3.29$ using the listed dimensions.
Characterization of inertial variation and aerodynamic forces
Unlike conventional aerial robots and fast-flapping insects, the AirPulse robot exhibits pronounced body undulation reminiscent of biological butterflies (Fig. 1D and movie S3). This motion originates from the large periodic variation of its inertial properties, which drives a strong coupling between the wing and body dynamics. Experimental measurements reveal that the fore-aft CG shifts by up to $\sim 3\%$ of body length, while the vertical displacement reaches approximately twice the body core height (Fig. 2A). Simultaneously, the principal pitch and yaw moments of inertia ($I_{yy}$ and $I_{zz}$) vary by as much as $\sim 2.5\times$ within a single wingstroke (Fig. 2B). According to the governing rotational dynamics of FWMAVs that follows from the angular-momentum balance,
where $\dot{I}_{yy}\dot{\theta}$ represents the coupling between time-varying inertia around body-$y$ axis and pitch rotation. In conventional aircrafts, $I_{yy}$ is nearly constant, so $\dot{I}_{yy}\dot{\theta}\approx 0$ and the body responds only to external torques ($\tau$). In the AirPulse robot, however, $I_{yy}(t)$ changes rapidly as the wings accelerate and decelerate, generating large internal reaction torques even without aerodynamic input. When $\dot{I}_{yy}<0$ (wings retracting from $70^{\circ}$ to $0^{\circ}$; Fig. 2B), rotational kinetic energy is drawn from the body to accelerate the wings; when $\dot{I}_{yy}>0$, energy is partially recovered. The alternating exchange of angular momentum between wings and body over successive strokes drives the observed body undulation. Scaling analyses further explain why the undulatory motion is more prominent at a smaller scale. Aerodynamic torque scales as $\tau_{\text{aero}}\propto l^{4}$, while inertial torque scales as $\tau_{\text{inertial}}\propto l^{3}$, yielding a ratio $\tau_{\text{inertial}}/\tau_{\text{aero}}\propto l^{-1}$ (where $l$ is the characteristic length). Therefore, inertial effects become increasingly dominant as system size decreases. The AirPulse robot operates near this transition regime, where wing inertia is a significant fraction of total body inertia, making $\dot{I}_{yy}\dot{\theta}$ non-negligible and leading to large body oscillations. In contrast, larger flapping robots, where wings contribute less to total inertia, exhibit aerodynamically dominated, nearly rigid-body motion.
To characterize the aerodynamic performance, static aerodynamic forces and moments (Fig. 2C) were first measured directly on the AirPulse robot mounted on a six-axis force-torque (F/T) sensor (BOTA MiniOne; 150 mN force resolution, 1.8 mNm torque resolution), with its CG aligned to the sensor center, installed in an open-circuit wind tunnel (1.8 m $\times$ 1.2 m $\times$ 1.0 m, turbulence intensity $7.5\%$, blockage ratio $5\%$) at a mean wind speed of 3 m/s. Data follow the convention of positive pitching moment as nose-up. Measurements at different flapping angles reveal a positive pitching-moment slope, confirming that the aerodynamic neutral point lies ahead of the CG. This longitudinal static instability enhances pitch responsiveness but implies that sustained flight cannot rely on passive aerodynamic stability and requires active control. Static lift measurements show a peak of 0.292 N at a $40^{\circ}$ flapping angle, exceeding the body weight of 0.26 N and allowing for gliding behavior. Nonetheless, static measurements cannot ensure sufficient lift during flapping flight maneuvers such as climbing or turning, where unsteady aerodynamic effects dominate.
To elucidate how flapping-wing morphology influences aerodynamic force generation, we then examined the aeroelastic behavior of butterfly-inspired wings characterized by broad planforms and vein-reinforced membranes. Such wings exhibit spatially heterogeneous stiffness arising from their intricate venation, which produces complex, phase-dependent deformations during flapping. This contrasts with the more uniform rod-reinforced wings commonly used in Diptera- or hummingbird-inspired robots [62, 63], where structural compliance is largely homogeneous. The venation-driven stiffness distribution in butterflies thus provides a natural mechanism for modulating local angle of attack and instantaneous force production throughout the wingstroke. To experimentally assess these effects, three wing variants were fabricated (Fig. 2E): (i) intact forewings with four longitudinal veins (R, M, Cu, A), (ii) intact forewings with two reduced reinforcement, and (iii) forewings with four veins paired with de-winged hindwings (hindwing membrane removed). The third variant isolates hindwing contributions, motivated by observations that hindwings primarily enhance maneuverability rather than generate essential lift [69].
Each wing variant was evaluated on a test platform (Fig. 2D) equipped with the aforementioned six-axis load cell along with a retroreflective-marker-based motion capture system (LUSTER FZMotion). Aerodynamic forces were obtained by subtracting inertial contributions (Materials and Methods). Results show that intact and de-winged four-vein wings produce similar in-cycle force profiles (Fig. 2, F and G), whereas two-vein wings exhibit distinct axial and vertical force patterns, indicating phase shifts caused by increased structural compliance. Furthermore, intact four-vein wings generate larger cycle-averaged axial and vertical forces than two-vein wings, but slightly lower than de-winged variants (Fig. 2H). This outcome likely reflects that the hindwing, in the current configuration, is neither biomimetically aligned with the forewing nor designed with a proper stiffness gradient. As a result, it flaps about an aerodynamically suboptimal axis, inducing additional drag and disrupting forewing wake flow. Wing deformation analysis reveals comparable flapping amplitudes and maximum bending angles (14.4° and 12.3°) near the downstroke horizontal position for intact four-vein and two-vein wings (Fig. 2, I and J). However, two-vein wings display a larger chordwise deflection of the forewing (Fig. 2K), confirming higher compliance in distal regions. These findings align with prior studies showing that trailing-edge flexibility reduces aerodynamic force in flexible wings with isotropic materials and simplified venation [62].
Figure 2: Inertial and aerodynamic characteristics. (A, B) Fore-aft and vertical CG displacement and principal inertia variation over a full wingstroke. (C) Static aerodynamic forces and moments under a 3 m/s horizontal wind; positive pitching moment is nose-up. (D) Test rig with six-axis force-torque (F/T) sensor and motion capture system for wing performance analyses. (E) Three wing variants: intact four-vein forewings (blue), two-vein forewings (orange), and four-vein forewings with de-winged hindwings (green). (F, G) Cycle-resolved axial and vertical aerodynamic forces for each wing variant. (H) Cycle-averaged axial and vertical forces for each wing variant. (I, J) Forewing bending angles during flapping for intact four-vein and two-vein wings. (K) Chordwise deformation illustrating higher distal compliance in two-vein wings.
Stroke timing asymmetry rhythm: a smooth and stable flapping pattern
Generating aerodynamic moments on small-scale flapping-wing robots requires low-dimensional modulation that is both numerically stable and kinematically smooth, particularly ensuring continuity of the first derivative of stroke angle, while maintaining mechanical simplicity. Biological fliers achieve this largely via subtle modulations of wingstroke kinematics [37]. In engineering practice, three principal kinematic strategies are commonly employed for generating periodic signals: stroke amplitude modulation (less common mechanism in real insect flight control), stroke angle offset modulation (shifting the neutral position, and therefore, the stroke plane; Fig. 3A), and stroke timing or velocity modulation (Fig. 3B). The first two approaches are straightforward and widely employed [22, 54, 34]. In contrast, stroke timing modulation generates asymmetric upstroke and downstroke velocities within a single wingbeat. This asymmetry produces a nonzero cycle-averaged drag force on each wing while maintaining stroke amplitude. Despite its biological relevance for rapid maneuvers such as body saccades, it has received comparatively few systematic studies [65, 67, 68, 14].
Existing stroke timing modulation schemes face a trade-off between stability and smoothness (Fig. 3, C and D). Polynomial phase-shaping methods yield differentiable kinematics but permit only limited up-down stroke asymmetry and become sensitive when parameters exit a narrow admissible range [65], which may result in erroneous flapping behavior. Split-cycle strategies allow larger asymmetry but introduce either instability when modulation exceeds feasible limits [67] or kinematic discontinuities at switching points [68, 14], which risk actuator saturation and structural excitation. These limitations motivate a stroke timing modulation law that ensures at least first-derivative continuity while expanding the range of attainable asymmetry in a stable manner.
where $\zeta$ is the flapping amplitude, $f$ is the nominal flapping frequency, and $A$ is the dimensionless modulation parameter for asymmetric stroke timing whose admissible range can be linearly normalized to $(0,1)$ for convenience. The cosine form of the modulation function $p(\cdot)$ is deliberately chosen to suppress phase offsets arising from initial errors (Supplementary Text). Formally, this formulation provides three guarantees (Materials and Methods): (i) the cycle-averaged frequency is exactly preserved for any constant $A$, allowing intra-cycle timing to be skewed without altering mean rate; (ii) stroke asymmetry scales linearly and monotonically with $A$, yielding a simple yet direct tuning rule; (iii) both stroke angle and its first derivative remain continuous, preventing abrupt accelerations that could saturate actuators or excite structural vibrations.
To maintain flight stability under time-varying control commands, the control system must transition between different asymmetry levels without inducing mechanical jars or phase lags between the wings. We achieve this by applying an Infinite Impulse Response (IIR) filter to the reciprocal variable $1/p$ rather than the modulation function $p$ itself (Fig. 3, H and I, and Materials and Methods). This nuance is essential for the time consistency of the rhythm because the instantaneous phase velocity is directly proportional to $1/p$, therefore smoothing this term ensures an unbiased mapping between the control input $A$ and phase increment (Fig. 3, J to L). In contrast, filtering $p$ would introduce a nonlinear lag that distorts the intended stroke timing and induces a cumulative phase drift. By processing the reciprocal variable, we preserve a linear relationship between the modulatory control demands and the resulting wing motion. This ensures that STAR remains both responsive to time-varying commands and immune to the timing errors that typically plague rhythmic systems, ultimately allowing for more predictable and effective attitude control.
Figure 3: Stroke timing modulation using the STAR generator. (A) Stroke angle offset modulation shifts the neutral stroke plane. (B) Stroke timing modulation produces asymmetric upstroke/downstroke velocities. (C, D) Limitations of existing polynomial phase-shaping and split-cycle methods. (EG) STAR generates continuous, monotonic up-down asymmetry with smooth first derivatives across the wingbeat cycle for varying modulation parameter AA. (H, I) Performance of IIR filtering applied to $1/p$. (JL) STAR with IIR filtering preserves cycle-averaged flapping frequency while enabling stable and linear tuning of stroke asymmetry without introducing phase distortion.
Mapping forces and torques from flapping kinematics
The flapping motion of biological butterflies is governed by Central Pattern Generators, neural circuits that generate rhythmic motor outputs without the need for rhythmic inputs. These networks exhibit strong adaptability, as their intrinsic neuronal properties and synaptic connections can be dynamically modulated through sensory feedback. Following this biological principle, we implemented an analytical nonlinear CPG model that integrates sinusoidal oscillator with STAR, described by
$y(t)=\zeta\sin(\omega(t))+\delta,$ (3)
where $\delta$ denotes the stroke angle offset and other variables are defined in (2). While Hopf oscillators are commonly used for robotic CPGs, they primarily ensure rhythmic stability during amplitude transitions with fixed phase velocity [71]. This property contrasts with insect flight, where differential stroke timing rather than amplitude modulation alone serves as the primary control mechanism. The sinusoidal oscillator with STAR (3) captures this mechanism directly, producing wing kinematics that more closely resemble those of real insects. The oscillator output drives two micro servos via PWM signals, allowing independent control of the left and right wings. The main modulatory parameters include flapping frequency $f$, amplitude $\zeta$, angle offset $\delta$, and the STAR parameter $A$, which determines stroke timing asymmetry.
A central challenge in designing flapping-wing robots is to determine control effectiveness. That is, how control inputs (or modulatory parameters) translate into measurable forces and torques, thereby enabling effective control allocation and accurate dynamical modeling. This challenge is particularly evident in the AirPulse robot, which operates with only two actuators in a 6-DOF workspace. Such underactuation poses a complex control-allocation problem, as limited inputs must generate coupled and nonlinear force-torque responses across multiple axes. Existing studies on butterfly-inspired robots [58, 59, 60, 61, 54] rarely provide systematic quantification of these relationships, leaving control effectiveness of such type of aerial robots insufficiently understood. To this end, we performed a series of experiments measuring the total flapping-induced forces and torques using the test rig (Fig. 2D) for establishing a quantitative foundation for the controller design of the AirPulse robot. While theoretical models exist such as quasi-steady ones based on instantaneous forces and kinematic parameters [40, 57, 44], unsteady mechanisms such as Leading-Edge Vortexs, rotational circulation, wake interference, and structural flexure render these first-principles modeling intractable for heterogeneous wing architectures [44]. Furthermore, the system’s strong sensitivity to kinematic variations, combined with unsteady aerodynamic effects, renders analytical control design overly sophisticated. The empirical approach therefore captures the complete set of physical effects, providing direct measurements that inherently incorporate aerodynamic-inertial coupling often neglected in simplified theoretical formulations.
Angle offset modulation exhibits more nuanced control authority. Symmetric increases in $\delta$ (both wings shifting upward; Fig. 3A) reduce total force while generating progressively larger pitch moments (Fig. 5F), resulting in nose-down torques consistent with free-flight observations. This confirms that symmetric angle offset modulation functions as a reliable mechanism for pitch control. In contrast, antisymmetric offset modulation (opposite shifts of the left and right wings) produces minimal changes in total force or yaw torque but induces pronounced roll moments when the left wing elevates and the right wing descends (Fig. 5G). Interestingly, this behavior seemingly contrasts with free-flight maneuvers, where the same configuration produces left turns. This indicates a fundamentally different turning mechanism from that of fixed-wing aircraft, in which roll moments generate banked turns. For the AirPulse robot, turning is likely initiated by adverse yaw, created when one wing experiences greater drag during the downstroke, yawing the robot’s nose toward that side. This yaw motion subsequently induces a sideslip in the opposite direction, which, together with the wing’s dihedral angle, produces a secondary rolling moment. Hence, the roll observed in flight is a consequence of yaw rather than its cause. This mechanism is empirically supported by free-flight observations, where roll excursions during turning remain small, indicating that the maneuver is primarily yaw-driven, similar to previously reported yaw turns in dragonflies [72]. The observed roll may even occasionally occur in the opposite direction (a slight right roll during a left turn) or lag behind the yaw motion. Overall, antisymmetric angle offset modulation proves to be an effective strategy for executing turning maneuvers.
Compared with previous modulation methods, stroke timing modulation via the STAR parameter $A$ exhibits distinct force-torque characteristics. Symmetric increases in $A$ (faster downstroke and slower upstroke; Fig. 3B) decrease mean pitch torque, generating a nose-up moment and establishing stroke timing as a complementary mechanism for pitch control alongside angle offset modulation (Fig. 5I). Antisymmetric modulation (asynchronous timing between wings) produces pronounced roll moments with minor yaw torque variations (Fig. 5J), consistently resulting in right-turn maneuvers confirmed in free-flight tests. Notably, STAR-based modulation shows greater variance in force and torque profiles than angle offset modulation, indicating phase-dependent alterations in periodic force generation (fig. 4) rather than amplitude-driven differences (fig. S3).
Figure 4: Forces and torques generated by stroke timing modulation. (A) Symmetric modulation (p_l_r_X): both wings have stroke timing value $X$. (B) Antisymmetric modulation (p_l_X_r_Y): left and right wings have stroke timing values $X$ and $Y$, respectively. Colors indicate different modulation settings.
These results highlight the complementary roles of angle offset and stroke timing modulation. Angle offset modulation primarily scales the magnitude of forces, whereas stroke timing modulation adjusts their phase within each wingstroke. The variance patterns support this distinction: STAR modulation alters the temporal phasing of instantaneous forces, while offset modulation changes their amplitude without affecting timing. These findings establish a quantitative mapping between modulation parameters and dynamic responses, forming a control effectiveness basis for hierarchical control architectures in which high-level commands modulate CPG parameters to achieve desired force-torque profiles.
Building on the established mapping between flapping modulation and force-torque generation, we developed a control architecture for free flight (fig. S4) composed of three modules: state estimator, attitude controller and CPG (Supplementary Text). The state estimator fuses raw IMU signals using a Madgwick filter, providing drift-compensated attitude estimates on a resource-constrained hardware in real time. These estimates are processed by an attitude controller based on PID, which computes errors from the desired references and outputs modulatory commands to the CPG. Lastly, CPG converts these commands into periodic flapping PWM signals, independently driving the two micro servos to achieve differential wing modulation between the left and right wings. In contrast to fast-flapping robots in the literature [11, 21, 29, 30], which can be effectively controlled using stroke-averaged models, the AirPulse robot exhibits strong dynamic coupling between its low-frequency flapping and body undulation (Fig. 1D and Fig. 6, F and H). This coupling violates the core assumption of timescale separation, rendering the conventional stroke-averaged approach invalid for our platform [73].
To achieve stable attitude regulation within this oscillatory regime, we employed an online estimation algorithm based on Recursive Least Squares (RLS) (Materials and Methods) that adaptively extracts the low-frequency mean of the pitch signal (Fig. 6, F and H). The estimated signal serves as a control-effective surrogate of the oscillatory pitch angle, enabling the computation of stable error and non-oscillatory control commands. Two complementary control pathways, namely angle offset and stroke timing modulation, were evaluated in free-flight experiments, demonstrating how distinct phase and amplitude control channels can be exploited to achieve climbing and turning maneuvers (Fig. 6, D and E; movies S4 and S5). Both modulation strategies achieved accurate pitch tracking by adjusting either the stroke offset (Fig. 6, F and G) or the stroke timing (Fig. 6, H and I), consistent with the force-torque relationships established in experimental characterization. Average power consumption remained low (5.88 W and 5.95 W, respectively), underscoring the energetic efficiency of the flapping mechanism. For directional control, yaw was selected as the primary variable because free-flight tests showed that turning arises from coupled roll-yaw dynamics, with roll excursions remaining minor. Stable turns were achieved using either modulation strategies (Fig. 6, J and L), with comparable power requirements (6.04 W and 5.87 W; Fig. 6, K and M). Compared to micro quadrotors [74], which draw approximately 10.6 W in hover, these results demonstrate that the AirPulse robot attains forward flight with substantially lower power demand, revealing the aerodynamic efficiency of dynamically coupled flapping-wing motion. These experiments validate a closed-loop control framework that integrates flapping kinematics, state estimation, and phase-coordinated modulation of wing motion to achieve stable undulatory free flight.
Discussion
The present study demonstrates the first and lightest untethered flight of a butterfly-inspired, two-winged tailless flapping-wing robot, AirPulse, in the peer-reviewed literature (Fig. 7A and Table 1), capable of executing climbing and turning maneuvers using only onboard sensing and control. The 26 g platform employs two micro servos to drive low-aspect-ratio wings with biomimetic venation, reproducing the heterogeneous stiffness distribution characteristic of real butterfly wings. Through systematic force-torque mapping, we identified how flapping parameters—including amplitude, frequency, stroke angle offset, and stroke timing—govern the generation of aerodynamic and inertial forces. Building upon these insights, we developed a flight controller that accounts for body undulation caused by cyclic variations in the inertial properties. This controller achieves stable pitch and yaw tracking via two complementary modulation strategies: angle offset control, and phase-domain rhythm generator (STAR) that enables smooth and stable modulation of stroke timing asymmetry.
The AirPulse robot demonstrates strong biological relevance to real butterfly flight, for instance, closely matching the swallowtail butterfly (Papilio thoas) in the ratio between aspect ratio ($\textrm{AR}=b^{2}/(2S)$, where $b$ is wingspan and $S$ is single wing area) and wing loading ($\textrm{WL}=mg/(2S)$, where $m$ is body mass) (Fig. 7B). Notably, its wing loading is among the lowest in previously reported butterfly-inspired robots [54, 66] (Table 1), suggesting improved aerodynamic efficiency, enhanced maneuverability, and the ability to sustain low-speed flight [50]. The aerodynamic regime of the AirPulse robot was quantified by its Reynolds number, calculated as $Re=\rho V\bar{c}/\mu\approx 5{,}600$, based on a mean aerodynamic chord of $\bar{c}=88.6$ mm and characteristic velocity $V=1.03$ m/s. This regime aligns with biological counterparts such as Vanessa cardui ($Re\approx 3{,}500$[51], confirming that the robot operates within the same viscous, low-Reynolds number region where unsteady mechanisms, including LEV formation, rotational lift, and wake capture, dominate aerodynamic force generation. In contrast, conventional rotary-wing robots of comparable scale typically operate at Re>25, [^0] Re>25{,}000, where airflow remains largely attached and steady. The AirPulse robot thus represents a fundamentally different aerodynamic paradigm that is highly reliant on vortex-dominated flow physics rather than quasi-steady lift generation.
To further elucidate the aerodynamic operating regime of the AirPulse robot, we analyzed the reduced frequency, defined as $k=\omega\bar{c}/(2U)$, where $\omega$ is the flapping angular velocity, $\bar{c}$ is the mean chord length, and $U$ is the forward flight speed. By inverting this relation, we estimated the corresponding forward velocity range for given $k$ values, thereby identifying the aerodynamic transition between quasi-steady ($k\ll 1$) and unsteady ($k>1$) regimes (Fig. 7C). The results further confirm that, for working flapping frequencies and chord dimensions, the AirPulse robot operates predominantly in the unsteady to transitional region, where vortex-dominated mechanisms play a central role in lift generation. Moreover, consistent with prior findings [48], aerodynamic efficiency increases with decreasing reduced frequency or non-dimensional wing loading, implying that the AirPulse robot not only reproduces key aerodynamic features of butterfly flight but also provides a controlled and measurable testbed for investigating the coupled inertial-aerodynamic mechanisms within a biologically relevant regime comparable to that of real butterflies [46, 52].
Despite these advances, several limitations and open challenges remain. The current PID-based controller, while effective in stabilizing pitch and yaw, requires empirical gain tuning and lacks sufficient robustness under dynamically varying aerodynamic loads or exogenous disturbances. The strong fluid-structure coupling intrinsic to flexible wings introduces additional nonlinearities that are not fully captured by classical control frameworks. Developing higher-fidelity dynamic models remains an active area of investigation, essential for the design of advanced controllers such as adaptive, model-based, or neural control architectures [75]. Furthermore, while our biomimetic venation design reflects spatial stiffness distribution, the influence of interspecific venation angles and vein topology on aerodynamic performance remains unresolved. Current datasets are insufficient to statistically determine whether observed angle variations across species represent evolutionary features or random morphological variance. Broader interdisciplinary efforts combining biological morphology, computational modeling, and robotic validation will be necessary to uncover such correlations. Our future work will also focus on integrating neuromorphic architectures to couple event-driven sensory modalities such as dynamic vision [76] and olfactory cues [77] for navigation in cluttered or GPS-denied environments. Incorporating such sensory-motor integration may ultimately allow flapping-wing robots to exhibit autonomous behaviors comparable to real insects, bridging the gap between biological and engineered flight systems.
@@ Acknowledgments
We are grateful to Prof. G.C.H.E. de Croon, Prof. Maurizio Porfiri, and Prof. Lu Fang for their insightful comments and feedback on this manuscript. W.G. acknowledges support from the China Postdoctoral Science Foundation under Grant Number 2025M781650. This work was also supported by Xinchen Qihang Inc.
@@
Corresponding authors: Weibin Gu, Chao Gao, Guyue Zhou
Email: chao.gao@cantab.net, {guweibin, zhouguyue}@air.tsinghua.edu.cn
@@
Wings were fabricated from Mylar film reinforced with carbon fiber spars (a 1.0 mm main spar, 0.8 mm leading edge and hindwing contour, and 0.5 mm longitudinal veins). The carbon reinforcement provided a high stiffness-to-weight ratio and enabled spatial stiffness gradients analogous to those found in butterfly wings. Alternative membranes were evaluated: TPU exhibited excessive compliance and poor adhesion, while silicone-coated polyamide (PA66) offered durability but imposed excessive mass. All tested membranes were flight-capable; however, PET was ultimately chosen for its balance of light weight, stiffness, tear resistance, adhesive compatibility, and optical transparency for motion capture. Wing fabrication involved laser-printing a venation template, overlaying the PET film, and bonding carbon rods using UV-curable adhesive and cyanoacrylate. Left-right wing pairs were mass-balanced within 0.05 g ($\leq 1\%$ of a single wing mass) to minimize asymmetries that could induce unbalanced aerodynamic forces or fuselage vibrations from uneven stroke timing or deformation [32].
@@
where $\{\mathbf{P}_{i}\}_{i=0}^{n-1}$ are the spline control points determined by smoothing least-squares fitting, and $N_{i,3}(u)$ are the cubic B-spline basis functions defined by the Cox-deBoor recursion [78] as
@@
Parameter Value Units
Sampled points ($N$) 477 points
Control points ($n$) 22 points
Data reduction 95.4%
RMS fitting error 0.672 pixels
Smoothing factor ($\alpha$) 0.5
Resampled points ($M$) 1000 points
Index $i$ $x_{i}$ (pixels) $y_{i}$ (pixels) Index $i$ $x_{i}$ (pixels) $y_{i}$ (pixels)
0 312.908971 -210.627073 11 176.467499 51.938866
1 256.030813 -211.484444 12 220.083362 52.401279
2 187.390279 -184.698024 13 263.699226 52.863692
3 95.897335 -120.190052 14 307.315090 53.326105
4 43.167102 -65.320731 15 338.611816 41.334592
5 4.851825 -16.166765 16 349.589403 11.000153
6 -2.349460 10.755896 17 340.247852 -29.334286
7 21.089190 29.768771 18 310.587162 -69.668725
8 63.968615 36.324384 19 260.607334 -110.003164
9 118.976443 44.915320 20 190.308367 -150.337603
10 147.655635 48.477453 21 99.690262 -190.672042
@@
Forewing images were collected primarily from published literature with documented clear venation patterns [79], supplemented by photographs of museum-grade specimens available from open-access sources (Wikipedia). Specimens were selected based on strict visibility criteria requiring a fully unobstructed forewing (dorsal or ventral) view with distinct Dc and Cu1 veins, which define key geometric axes for comparative analysis. The dense and pigmented scales on butterfly wings often obscure vein boundaries, substantially constraining the number of usable samples. Consequently, the final dataset included 24 different specimens representing three families, namely Nymphalidae ($n=8$), Pieridae ($n=8$), and Papilionidae ($n=8$). For each specimen, the forewing image was first converted to grayscale and manually annotated with yellow solid lines indicating the Dc and Cu1 veins (fig. S2). The annotated regions were then segmented in HSV color space to extract the obtuse inter-vein Dc-Cu1 angle. Future morphological studies could expand statistical results by employing chemical bleaching or depigmentation techniques to enhance venation contrast and enable automated image segmentation.
@@
Wing performance was assessed using a custom flapping apparatus (Fig. 2D). Retroreflective stickers were selected for their low mass, which is the lightest available option despite still adding some weight. Increasing the number or size of markers could improve tracking accuracy but would also alter wing dynamics due to extra weight.
@@
Each wing variant underwent same flapping cycles while synchronized force and kinematic data were recorded at 222 Hz. The raw force signals were processed through a computational pipeline that first identified active flapping periods by detecting servo motor activation thresholds in the PWM command signals. Stroke reversal points were determined via second-order differentiation of positional data to precisely segment individual flapping cycles. Aerodynamic forces were isolated by subtracting inertial contributions using a phase-locked averaging approach. Specifically, a perforated counterpart was fabricated for each wing variant to approximate the original mass and inertial properties while substantially reducing aerodynamic loading. Forces measured under identical flapping conditions with these perforated wings were treated as inertial baselines and subtracted from the corresponding intact-wing measurements. For each wing configuration, 40 consecutive flapping cycles were selected from steady-state operation and processed with a fourth-order Butterworth low-pass filter (50 Hz cutoff) to remove high-frequency noise. The filtered force data were temporally normalized to the interval $[0,1]$ to enable cycle-wise comparison. Component-wise force signals ($F_{x}$, $F_{y}$, $F_{z}$) were then averaged and their variances computed to characterize aerodynamic force profiles across the normalized flapping period. Processed forces were further analyzed via polar coordinate transformation, where the angular coordinate represented normalized time ($\theta_{\mathrm{polar}}=2\pi t_{\mathrm{norm}}$) and the radial coordinate denoted force magnitude with an offset ensuring positivity (Fig. 2, F and G). Time-integrated forces were computed using trapezoidal numerical integration for both signed and absolute components, providing quantitative measures of net impulse and cumulative force production. This analysis enabled comprehensive comparison of force magnitude, temporal phasing, and generation patterns across different wing morphological variants.
@@
Wing deformation was quantified by projecting three-dimensional marker positions onto a plane orthogonal to the wing rotation axis, isolating bending motion from the overall flapping trajectory. Bending angles were defined as the angular deviation between proximal and distal segments, with maxima identified separately for downstroke and upstroke phases (Fig. 2, I and J). Hindwing kinematics were characterized by computing elevation angles relative to the wing hinge throughout each flapping cycle. Forewing kinematics were derived from markers placed along the anterior (leading edge) and posterior (trailing edge) margins of each wing section, enabling direct quantification of instantaneous wing orientation relative to the airflow. Discrete marker data were interpolated to generate continuous angle-time curves, enabling phase-resolved comparison of fore-/hind-wing kinematics across wing variants.